OpenStudy (anonymous):

Find the rate of change between the two points: (11, 250) and (16, 200) where x is in days and y is in dollars.

OpenStudy (alexandervonhumboldt2):

use slope formula again

OpenStudy (alexandervonhumboldt2):

use slope formula again\[\frac{ y_1-y_2 }{ x_1-x_2}\]

OpenStudy (anonymous):

Rate of change \[\frac{\Delta y}{\Delta x}\]

OpenStudy (anonymous):

You have it reversed @AlexandervonHumboldt2

OpenStudy (anonymous):

is it -10

OpenStudy (alexandervonhumboldt2):

yeah -10.

OpenStudy (rhr12):

but its about day and dollars. How it can be a negative 10?

OpenStudy (alexandervonhumboldt2):

yeah....hmmmm

OpenStudy (alexandervonhumboldt2):

some idiot made the question

jimthompson5910 (jim_thompson5910):

the negative 10 means that the person is losing $10 a day

OpenStudy (alexandervonhumboldt2):

yeah!

OpenStudy (alexandervonhumboldt2):

awesome work jim!

OpenStudy (anonymous):

You have the right answer, but the equation is flopped... Needs to be \[\huge \frac{ y_2-y_1 }{ x_2-x_1}\] @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

it doesnt matter

jimthompson5910 (jim_thompson5910):

@CShrix it turns out that it doesn't matter if you do y2-y1 or y1-y2. The only thing that matters is if you keep the order the same If you did y2-y1, then you have to do x2-x1 If you did y1-y2, then you have to do x1-x2